### v1

@@ 4 Study Findings
In our study, we examined the effects of four common data bugs: label noise, class imbalance, concept drift, and missing preprocessing. To address each research question, we constructed and trained a comprehensive dataset of 120 buggy models, with 30 models dedicated to each of the four bug types. Additionally, we trained 30 models on clean data to serve as a baseline, helping us establish the expected training behaviour. We used the W&B logging framework wandb to capture the training behaviours of both faulty and clean models. This section presents our findings, focusing on the most common symptoms of these data bugs across three data types: code-based, text-based, and metric-based.
@@ 4.1.1 Impact of Data Quality
(a) Inconsistent Learning: Inconsistent learning in neural networks refers to the phenomenon where different layers or components of a network learn at varying rates or effectiveness, leading to suboptimal overall performance bengio1994learning . We observe inconsistent learning across the neural network layers when models were trained on low-quality data. As shown in Table. 4, 80% of the models trained on data with label noise and 56.67% of the models trained on data with concept drift demonstrated inconsistent learning. Our manual analysis of these models shows that $\approx$20of their layers were affected, especially the attention and output layers. Due to the data bugs, these layers learn very slowly, as shown by their near-zero biases ($<$ 0.01) neal2018modern . In contrast, the layers unaffected by data quality issues learn quickly (bias >=>=> = 0.5). Attention and the output layers of the models also showed normal bias values (bias >=>=> = 0.5). Thus, the disparity in the bias values across the different layers of a neural network trained on low-quality code data results in inconsistent learning.
@@
(b) Reduced Model Capacity: We found that models trained on low-quality data exhibited a reduced capacity to capture complex patterns from data. As shown in Table 4, 86.67% of models trained on data with label noise and 63.33% of models trained with concept drift demonstrated reduced capacity. Our analysis of these models revealed that 90-95% of their layers were affected, particularly embedding, attention, and dense layers. These layers exhibited notably small weights (e.g., $\mu\approx 0.011$, $\mu\approx 0.047$ for models trained with label noise and concept drift, respectively). These values were significantly lower than the typical weight values of 0.1 to 0.3, which were observed when the models were trained on clean data. The small and tightly clustered weights suggest a limited range in the representational power of a model’s neurons, which indicates the reduced capacity of the corresponding model liu-etal-2020-understanding ; bengio1994learning .
(c) Gradient Instability: We also noted gradient instability when models were trained on low-quality data. As shown in Table 4, 73.33% of models trained on data with label noise and 53.33% with class imbalance exhibited gradient instability. Our analysis of these models revealed that 30-40% of their layers were affected, especially the attention and dense layers. These layers demonstrated extreme gradient values, ranging from $1\times 10^{-9}$ to $1\times 10^{1}$, which was a significant deviation from the range of $1\times 10^{-3}$ to $1\times 10^{-1}$, as observed in the baseline models trained on clean data. This wide range of gradient magnitudes leads to unstable learning, where some layers update their weights rapidly, and others remain almost static. This gradient instability results in suboptimal model performance, as the model fails to develop a coherent understanding of code data across all of its layers pascanu2013difficulty .
@@
Data Quality Issues Incosistent Learning Reduced Model Capacity Gradient Instability
@@ 4.1.2 Impact of Missing Preprocessing Operations
(a) Slow Convergence: In code-based data, preprocessing operations such as tokenization and stopword removal are crucial, as they enable the accurate parsing and analysis of source code. Our analysis reveals that 66.67% of the models trained with missing preprocessing operations had slower convergence. We also found that $\approx$15of the layers in these models suffered from slow convergence, especially the output layers. The slow convergence of the parameters of the output layers indicates that no class was probable enough, which affects the decision-making ability of a model. Their weights and gradients were close to zero ($<$ $10^{-6}$) as opposed to $1\times 10^{-3}$ to $1\times 10^{-1}$ from the baseline models burden1997numerical . Such low gradients and weights often result in inefficient gradient flow, which affects a model’s ability to converge towards the optimal parameter values.
(b) Incorrect Learning: We noted that the lack of preprocessing operations in code-based data led to inconsistent or incorrect learning. Our analysis showed that the models trained on clean datasets exhibited a uniform and bounded distribution in their bias values (e.g., ranging from 0.34 to 0.72). Such a distribution enables better learning and generalization to unseen data srivastava2014dropout . On the contrary, we noted unbounded bias values in the models trained on data without preprocessing. Some layers displayed bias values near zero ($\approx$0), while others had bias values exceeding 1.0. This behavior was prevalent in 53.33% of the models trained without preprocessing, affecting $\approx$60of their layers. These inconsistent bias patterns indicate that the models struggle to leverage biases effectively, resulting in a limited set of learned weights primarily clustered near the origin ($\approx$0). Consequently, this leads to poor generalization and reduced “guessing power”, degrading the model’s ability to comprehend and analyze source code effectively Bennett_2017 .
(c) Distorted Distributions: Missing preprocessing steps can significantly distort the distribution of weights from deep learning models, as suggested by our analysis. We found that 70% of the models trained without preprocessing were affected. Our analysis of these models showed that 80% of their layers exhibit distorted distributions, particularly in embedding, attention, and dense layers. They demonstrated a skewness of $>1$ and a kurtosis of $|k|>>>3$, which are significantly higher than the skewness (0.4) and kurtosis (1.3) of the model weights trained on clean data. Such high skewness and extreme kurtosis indicate that the weights and gradients of an affected model are concentrated at a point farther right from the baseline counterparts. That is, small changes in input can cause a large change in output, which makes the model highly sensitive to input changes, reducing its robustness and reliability.
@@
4.1.3 Impact analysis of Label Noise using Attention Mechanisms
@@
Attention weights have been effectively used in software engineering research to understand the correlation between the input and a model’s predictions (outputs) cao2024systematic . In our study, we also analyze the attention weights of a model to understand the impact of data quality issues on the model’s training. Our study examined various data quality issues, and we found that label noise had the strongest impact on models trained with code-based data (see Table 4). Based on this finding, we decided to conduct a qualitative analysis to gain a deeper understanding of how label noise affects code-based data. To investigate this issue, we trained 60 models in total. We divided these models into four groups of 15 models each: (a) models trained on a clean Devign dataset, (b) models trained on a clean BigVul dataset, (c) models trained on a Devign dataset with label noise, and (d) models trained on a BigVul dataset with label noise. For each group, we calculated the average attention weights across all 15 models. This approach helps us identify consistent patterns and mitigate the impact of non-determinism in deep learning. We discuss our findings below.
@@
Bug-free Data: Models trained on bug-free datasets on vulnerability detection demonstrate a strong focus on tokens critical for identifying potential security flaws. This focus is evident in the BigVul dataset, which contains source code from Chrome, Linux, Android, and Tcpdump. When analyzing models trained on this dataset, we observe high attention weights on key functions and data structures related to security vulnerabilities. For instance, rds6_inc_info_copy (87.3796) and struct rds_incoming (64.1523) receive significant attention, indicating their relevance to network-related vulnerabilities. The models also heavily emphasize network fields such as laddr (85.7431) and faddr (86.2221), demonstrating their awareness of potential security issues in network communications. The Devign dataset contains source code from Linux, FFmpeg, Qemu, and Wireshark. Models trained on this dataset focus on memory-related operations and type conversions, which are one of the primary causes of security vulnerabilities Stevens_Fenner_Rudoff_2013 . Tokens like uint16_t (88.8645) and uint64_t (86.9365) receive very high attention weights, which reflects their importance to proper data handling. Critical function calls for cross-platform compatibility, such as le16_to_cpu (82.2068) and le64_to_cpu (80.8885), are also strongly emphasized. Models trained on both datasets show significant attention to error-checking patterns and memory operations like sizeof (30.5034) and if(!conn) (32.1079). These attention weights suggest the models have developed a comprehensive understanding of code elements related to security flaws from the datasets.
@@
Buggy Data: In contrast, the models trained on buggy data showed their inability to focus on tokens crucial for vulnerability detection. In BigVul, the model assigns significant weights to non-vulnerable tokens or variables such as int (37.6882), rt (36.5832), and nl (41.7406). Similarly, in Devign, the model focuses its attention on various elements like static (50.7911), CCIDBus (55.7003), and Device (43.4922). In other words, the model fails to prioritize the tokens that might signal potential vulnerabilities, such as unsafe operations or improper error handling. Moreover, the models trained on data with label noise pay considerable attention to non-informative tokens, such as newline characters (26.1465 in BigVul, 54.1878 in Devign) and empty tokens (8.2194 in BigVul, 10.1193 in Devign). This diffused attention pattern, evident across both datasets, suggests that these models struggle to distinguish between benign code patterns and those that may introduce vulnerabilities. This limits their effectiveness in identifying security-critical code sections.
@@
Summary of RQ1: Models trained on code-based datasets containing data bugs exhibit many issues, including inconsistent learning across layers, reduced model capacity and gradient instability. Furthermore, missing preprocessing can lead to slow convergence, incorrect learning patterns, and distorted distributions, impairing the models’ ability to understand and process code effectively.
@@
4.2.1 Impact of Data Quality Issues
(a) Abnormal Weight Distribution: We observed abnormal layer weights in the models trained on low-quality text data. As shown in Table 5, 93.33% of the models trained on data with concept drift exhibited this phenomenon. Our analysis of these models revealed that 80-85% of their layers were affected, particularly the convolutional and dense layers. These layers demonstrated unusually high weight variances (e.g., $\sigma^{2}\approx 2.3$), significantly deviating from their baseline counterparts (e.g., mean weight variance $\sigma^{2}\approx 0.64$). The observed abnormal weight distribution can be attributed to the dynamic nature of text data in software engineering. As software and technology evolve, the corresponding text-based data also changes, which causes shifts in the underlying data distribution. This evolution in software and text can affect how the text is represented, and the relevant weights and biases are learned by the models. As a consequence, older and newer data points might require different sets of optimal parameters. During model training, the text data with concept drift generate large gradient signals to push the parameter values towards the optimal settings for their respective time periods, which could have led to abnormal weight distribution in the models.
(b) Gradient Skewness: We also observed significant gradient skewness in the models trained on low-quality text data. As shown in Table 5, 76.67% of the models trained on data with concept drift and 56.67% with label noise exhibited this phenomenon. Our analysis revealed that 80-90% of the layers in these models were affected, particularly the convolutional layers and batch normalization layers. The gradient distribution of these layers demonstrated substantial skewness ($\gamma\approx 2.9$ for models trained with concept drift, and $\gamma\approx 1.2$ for models trained with label noise), far exceeding the baseline skewness values ($\gamma\approx 0.2$). As discussed above, concept drift in text-based data can lead to large gradient signals during training. This not only affects weight distribution but also could make the gradients skewed. Large gradient signals shift the overall distribution towards higher values, causing positive skewness. Similarly, label noise can also cause gradient skewness. In our datasets with label noise, similar data points have different labels due to incorrect labelling. When a model encounters these mislabeled examples during training, it tries to learn a decision boundary that separates these similar points with different labels, which is theoretically impossible. This leads to large prediction errors, as the model’s output significantly differs from the (incorrectly assigned) target labels. These large errors, in turn, result in large gradient updates during backpropagation. As the training process continues, these inconsistent and large gradient updates accumulate, causing the overall gradient distribution to be skewed.
(c) Overfitting: We noticed overfitting in the models trained on low-quality text data. As shown in Table 5, 86.66% of models trained on data with class imbalance exhibited overfitting. Our analysis revealed that 60-65% of the layers in these models were affected, particularly the dense layers. The median bias value of these layers ($med\approx 1.1$) is much higher than their baseline counterparts ($med\approx 0.07$). When a dataset has a disproportionate number of samples in one class compared to the others, the model tends to become biased towards the majority class. As a result, the model’s biases become skewed towards the majority class, which results in the model overfitting the training data.

### v2

@@ 4 Study Findings
In our study, we examined the effects of four common data bugs: label noise, class imbalance, concept drift, and missing preprocessing. To address each research question, we constructed and trained a comprehensive dataset of 120 buggy models, with 30 models dedicated to each bug type. Additionally, we trained 30 models on clean data to serve as a baseline, helping us establish the expected training behaviour. We used the W&B logging framework wandb to capture the training behaviours of both faulty and clean models. This section presents our findings, focusing on the most common symptoms of these data bugs across three data types: code-based, text-based, and metric-based.
@@ 4.1.1 Impact of Data Quality
(a) Near-Zero Biases: When we trained our models using buggy data, we observed near-zero biases in several layers. Normal learning is characterized by well-defined bias values (e.g., $\geq$0.5), while near-zero biases (e.g., $<$ 0.01) might suggest learning difficulties neal2018modern . The bias parameter introduces non-linearity in the network and contributes to the learning process zhang2024bias ; williams2024expressivity . Thus, near-zero bias values might indicate a lack of non-linearity in the neural network. As shown in Table 4, 76.67% of the models trained on data with label noise and 50.00% of the models trained on data with concept drift demonstrated near-zero biases. Our manual analysis shows that $\approx$20of their layers were affected, especially the attention and output layers, which exhibited near-zero biases ($<$ 0.01) due to data bugs. In contrast, unaffected layers (e.g., embedding layer) maintained normal bias values ($\geq$ 0.5) and had regular parameter updates during backpropagation. These disparities in bias values across different layers indicate inconsistency in the learning capabilities of the layers ye2017importance . We also observed performance deterioration in the models with near-zero biases. As shown in Table 5, LineVul witnessed an 8.12%-8.39% decline across various performance metrics. CodeBERT exhibited similar behaviours for data bugs, with metrics decreasing by $\approx$6%.
@@
(b) Smaller Weights: We observed small weights in our models trained on buggy data. According to existing studies, small and tightly clustered weights in a neural network indicate reduced representational power of the neurons liu-etal-2020-understanding . As shown in Table 4, 83.33% of models trained on data with label noise and 56.67% of models trained with concept drift demonstrated a symptom of very small weights. Our analysis of these models revealed that 90-95% of their layers were affected, particularly embedding, attention, and dense layers. These layers exhibited small weights (e.g., $\mu\approx 0.011$, $\sigma\approx 0.009$ and $\mu\approx 0.047$, $\sigma\approx 0.038$), significantly lower than those observed in models trained on clean data (e.g., $\mu\approx 0.23$, $\sigma\approx 0.11$). These small weights had a significant impact on the model’s performance, causing a 14.77%-16.48% and 6.82%-13.17% reduction in various evaluation metrics for our baseline techniques (LineVul, CodeBERT).
(c) Gradient Instability: When we trained our models using buggy data, we observed instability in their gradients. From the perspective of neural networks, gradient instability indicates extreme variations in model parameter updates, which prevents effective signal propagation during the backpropagation step storm2020unstable . Moreover, existing literature has shown that the extreme gradient values might lead to exploding gradients and vanishing gradients pascanu2013difficulty . We found that models trained on buggy data exhibited significant gradient instability. As shown in Table 4, 68.33% of models trained on data with label noise and 53.33% with class imbalance demonstrated this issue. Our analysis revealed that 30-40% of their layers were affected, especially the attention and dense layers. We found that these layers have a gradient range between $10^{-9}$ and $10^{1}$, which is significantly larger than the gradient range from models trained on clean data — ($10^{-3}$ to $10^{1}$). These gradient instabilities also had a notable impact on the model’s performance, causing an 11.75%-13.01% and 10.01%-11.16% reduction in various performance metrics for our two baseline techniques (LineVul, CodeBERT).
@@
Data Quality Issues Near-Zero Biases Smaller Weights Gradient Instability
LineVul
@@
CodeBERT
Label Noise 73.33% 80.00% 63.33%
Class Imbalance 50.00% 60.00% 53.33%
Concept Drift 43.33% 50.00% 26.67%
Table 5: Impact of Data Quality Issues on Model’s Performance for Code-Based Data
Symptoms Accuracy Precision Recall F1-Score
LineVul
Performance on Clean Data 93.62% 94.24% 83.12% 88.45%
Near-Zero Biases 85.23% 86.12% 74.89% 80.13%
Smaller Weights 77.14% 78.92% 68.35% 73.24%
Gradient Instability 81.87% 81.23% 71.96% 76.31%
CodeBERT
Performance on Clean Data 67.25% 61.25% 58.63% 55.25%
Near-Zero Biases 59.87% 56.12% 50.23% 52.94%
Smaller Weights 54.13% 49.26% 47.89% 48.43%
Gradient Instability 57.24% 52.87% 49.12% 50.86%
@@ 4.1.2 Impact of Missing Preprocessing Operations
(a) Slow Convergence: When we trained our models without any appropriate preprocessing operations, they converged slowly. That is, these models took longer than expected to reach optimal parameter values during training. To identify this issue, we analyze the time (i.e., epochs) required for convergence. Out of all the studied preprocessing operations for code-based data, missing line separation had the most severe impact, affecting 70.00% of LineVul and 76.67% of CodeBERT models. The models trained with proper preprocessing operations converged to the optimal parameters in $\approx$3.7 epochs, whereas the models without such preprocessing could not converge even after 5 epochs. Furthermore, the impact of slow convergence on the model’s performance was evident, causing a 10.01%-11.14% and 3.11%-8.40% reduction in various performance metrics for our two baseline techniques (LineVul, CodeBERT).
(b) Extreme Bias Values: When we trained the models without proper preprocessing, we observed a high range in their bias values. The absence of stop word removal had the highest impact, affecting 56.67% of LineVul and 63.33% of CodeBERT models. Our analysis revealed that $\approx$60of their layers were affected, showing widely dispersed bias values ranging between 0.03 and 1.97, significantly different from those of error-free models: (e.g., 0.34 to 0.72). Contained bias values indicate stable parameter distribution and learning patterns during model training williams2024expressivity . We also found that models with extreme bias values suffer from a substantial performance degradation, causing a 13.77%-17.32% and 7.72%-13.12% reduction in various performance metrics for our two baseline techniques (LineVul, CodeBERT).
(c) Skewed Parameter Distributions: When we trained our models without appropriate preprocessing, we also observed skewed weight distributions across the model’s layers. The absence of sequence length normalization had the highest impact, affecting 73.33% of LineVul and 83.33% of CodeBERT models. Our analysis revealed that 80% of their layers were affected, particularly embedding, attention, and dense layers. In models trained on clean code data, weights exhibited moderate skewness (0.4) and kurtosis (1.3), indicating balanced distributions. However, models trained without preprocessing showed high skewness ($>1$) and extreme kurtosis ($|k|\gg 3$), deviating significantly from clean data distributions. We also noticed a performance degradation in the affected models, causing a 10.16%-13.01% and 5.39%-10.51% reduction in various performance metrics for our two baseline techniques (LineVul, CodeBERT).
Table 6: Manifestations of Missing Preprocessing in Code-Based Models
Preprocessing Issues Slow Convergence Extreme Bias Values Skewed Parameter Distributions
LineVul
Stop Word Removal 68.33% 56.67% 70.00%
Line Separation 70.00% 50.00% 66.67%
Sequence Length Normalization 63.33% 53.33% 73.33%
CodeBERT
Stop Word Removal 73.33% 63.33% 76.67%
Line Separation 76.67% 60.00% 70.00%
Sequence Length Normalization 70.00% 56.67% 83.33%
Table 7: Impact of Missing Preprocessing on Model’s Performance for Code-Based Data
Symptoms Accuracy Precision Recall F1-Score
LineVul
Performance on Clean Data 93.62% 94.24% 83.12% 88.45%
Slow Convergence 83.14% 84.23% 71.98% 77.62%
Extreme Bias Values 77.14% 76.92% 69.35% 72.89%
Skewed Parameter Distributions 81.87% 81.23% 72.96% 76.85%
CodeBERT
Performance on Clean Data 67.25% 61.25% 58.63% 55.25%
Slow Convergence 58.87% 54.12% 50.23% 52.14%
Extreme Bias Values 54.13% 48.26% 46.89% 47.53%
Skewed Parameter Distributions 57.24% 51.87% 48.12% 49.86%
@@
4.1.3 Impact analysis of Label Noise in Training Data using Attention Mechanisms
@@
Attention weights have been used in software engineering research to analyze the correlation between the input and the output of attention-based models cao2024systematic ; wang2020detecting ; ma2024unveiling . In our study, we also analyze the attention weights of our model to understand the impact of data quality issues on the model’s training. Among four data quality issues, we found that label noise had the strongest impact on models trained with code-based data (see Table 4). Thus, we conduct a qualitative analysis to better understand how label noise affects a model during training. Towards this, we trained 60 models in total. We divided these models into four groups of 15 models each: (a) models trained on a clean Devign dataset, (b) models trained on a clean BigVul dataset, (c) models trained on a Devign dataset with label noise, and (d) models trained on a BigVul dataset with label noise. For each group, we calculated the average attention weights for the tokens in the BigVul dataset across all 15 models. By analyzing 15 models from each group, we aimed to mitigate the impact of non-determinism in deep learning. We discuss our findings below.
@@
Bug-free Data: Models trained on the bug-free versions of our datasets demonstrate a strong focus on the tokens related to security flaws. BigVul dataset contains source code from Chrome, Linux, Android, and Tcpdump. When analyzing models trained on BigVul, we observe high attention weights for key functions and data structures related to security vulnerabilities. For instance, rds6_inc_info_copy (87.3796) and struct rds_incoming (64.1523) receive significant attention, indicating their relevance to network-related vulnerabilities. The models also heavily emphasize network fields such as laddr (85.7431) and faddr (86.2221), demonstrating their awareness of potential security issues in network communications. The Devign dataset contains source code from Linux, FFmpeg, Qemu, and Wireshark. Models trained on this dataset focus on memory-related operations and type conversions, major causes of security vulnerabilities Stevens_Fenner_Rudoff_2013 . Tokens like uint16_t (88.8645) and uint64_t (86.9365) receive very high attention weights, which reflects their importance to data handling. Critical function calls for cross-platform compatibility, such as le16_to_cpu (82.2068) and le64_to_cpu (80.8885), are also strongly emphasized. Models trained on BigVul and Devign separately show significant attention to error-checking patterns and memory operations like sizeof (30.5034) and if(!conn) (32.1079). These attention weights suggest the models have developed a comprehensive understanding of code elements related to security flaws when the datasets do not contain any label noise.
(a) Attention weights assigned to salient tokens (i.e., related to security vulnerability) in the BigVul dataset by bug-free and buggy models.
(b) Attention weights assigned to non-vulnerable tokens in the BigVul dataset by bug-free and buggy models.
Figure 2: Comparison of attention weights between bug-free and buggy models for both important and non-important tokens. The bug-free model (green) shows higher attention weights for important tokens and lower weights for non-important tokens, while the buggy model (red) shows the opposite pattern.
@@
Buggy Data: In contrast, our models trained on the datasets containing label noise showed their inability to focus on appropriate tokens. For example, our models trained on BigVul (with label noise) assign significant weights to such tokens that might not be related to security vulnerability: int (37.6882), rt (36.5832), and nl (41.7406). Similarly, the models trained on LineVul (with label noise) focus their attention on various elements like static (50.7911), CCIDBus (55.7003), and Device (43.4922). In other words, the models fail to prioritize the tokens that might signal potential vulnerabilities, such as unsafe operations or improper error handling. Moreover, these models assign big weights to non-informative tokens, such as newline characters (26.1465 in BigVul, 54.1878 in Devign) and empty tokens (8.2194 in BigVul, 10.1193 in Devign). This diffused attention pattern, evident across both datasets, suggests the models have difficulty separating safe code from potentially vulnerable code. This limits their effectiveness in identifying security-critical code sections.
We also compare the attention values for individual tokens across the models. Figure 2 shows fundamentally different attention patterns between bug-free and buggy models. The bug-free models trained on the BigVul dataset exhibit strong discrimination power by assigning high weights (80-89) to security-critical tokens like uint16_t (88.9) and network identifiers (laddr, faddr) (85-87), while maintaining minimal attention (1-8) on non-essential elements. In stark contrast, the buggy models pay little attention (15-40) to security-critical tokens while heavily focusing on basic syntax elements and identifiers - CCIDBus (55.7), static (50.8), and newline characters (54.2). Given the weights to appropriate tokens, bug-free models have a higher chance of differentiating between vulnerable and benign code, where the model with a noisy dataset could fall short. Our experimental results (e.g., Table 5) also show strong evidence to support this.
@@
Summary of RQ1: Models trained on code-based datasets containing data bugs exhibit many issues, including near-zero biases, smaller weights, and gradient instability. Furthermore, missing preprocessing can lead to slow convergence, extreme bias values, and skewed parameter distributions, impairing the models’ ability to understand and process code effectively.
@@
4.2.1 Impact of Data Quality
(a) Abnormal Weight Distribution: When we trained our models using buggy text data, we observed abnormal weight distributions across multiple layers. According to existing studies, abnormal weight distributions can reduce a model’s effectiveness in capturing important features liu-etal-2020-understanding ; ye2017importance . As shown in Table 8, 81.67% of the models trained on data with concept drift demonstrated abnormal weight distributions. Our manual analysis shows that 80-85% of their layers were affected, especially convolutional and dense layers. The models trained on the buggy text exhibited high weight variances (e.g., $\sigma^{2}\approx 2.3$) significantly deviating from baseline values (e.g., $\sigma^{2}\approx 0.64$). We also observed notable performance deterioration in models with abnormal weight distributions. As shown in Table 9, DCCNN witnessed a 5.18%-7.28% decline across various performance metrics. CodeBERT+MLP exhibited similar behavior, with performance metrics decreasing by 3.50%-4.10%.
(b) Gradient Skewness: We also observed significant gradient skewness in our models trained on buggy data. Gradient skewness indicates an asymmetry in the gradient distribution curve, with a heavy tail on any side. According to existing studies, asymmetric gradient distributions can lead to exploding and vanishing gradients, causing unstable parameter updates pascanu2013difficulty . As shown in Table 8, 68.33% of models trained on data with concept drift and 48.33% with label noise demonstrated gradient skewness. Our analysis of these models revealed that 80-90% of their layers were affected, particularly convolutional and batch normalization layers. These layers had the gradients with significantly higher skewness values (e.g., $\gamma\approx 2.9$ for concept drift), significantly higher than those observed in models trained on clean data (e.g., $\gamma\approx 0.2$). As shown in Table 9, these skewed gradients also had a substantial impact on the model’s performance, causing a 7.49%-9.44% and 5.66%-6.50% decline in various performance metrics for our baseline techniques (DCCNN, CodeBERT+MLP).
(c) Overfitting: When we trained our models using buggy text data, we observed overfitting, which can be characterized by unusually high bias values. From the perspective of neural networks, overfitting occurs when models memorize training data patterns instead of learning the patterns williams2024expressivity ; zhang2024bias . Unlike weights, bias terms are added directly to the layer outputs without normalization. When these bias values are abnormally high (e.g., $med\approx 1.1$ for models trained on buggy data versus the baseline models’ values of $med\approx 0.07$), they effectively dominate the computation, and this reduces the impact of the weights and inputs on the final prediction. Because of such high bias values, the neurons consistently produce similar activation values irrespective of the inputs, which causes the model to essentially memorize fixed responses. We found that models trained on imbalanced data exhibited significant overfitting in terms of extreme bias values. As shown in Table 8, 85.00% of models trained with class imbalance demonstrated this issue. Our analysis revealed that 60-65% of their layers were affected, especially the dense layers. We found that these layers have bias values around $med\approx 1.1$, significantly larger than the bias values from the baseline models – $med\approx 0.07$. Overfitting also had a substantial impact on the model’s performance, as DCCNN’s and CodeBERT+MLP’s performance metrics dropped by 8.05%-9.26% and 5.50%-7.95%, respectively (Table 9).